Your grandmother will be gifting you $150 at the end of each month for four years while you attend college. At a discount rate of 3.7 percent, what are these payments worth to you on the day you enter college?
Answer #1

These payments are worth to you on the day you enter college $ 6,682.99

Present value = Monthly payment * Present value of annuity of 1
= $           150 * 44.55324
= $ 6,682.99
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= 44.553239 i = 3.7%/12 = 0.003083
n = 4*12 = 48